Given the function K(s)=2s+9, find each of the following.
K(4)=
K (-7)=
K (6.6)=

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

Given the function K(s)=2s+9,
find each of the following.
K(4)= 2*4+9 = 17
-------------------
K(-7)=2(-7)+9 = -14+9 = -5
-------------------
K (6.6)= 2(6.6)+9 = 13.2+9 = 22.2
===================================
Cheers,
Stan H.
=====================

Answer by rfer(16322)

k(4)=2*4+9
k(4)=8+9
k(4)=17